Use the following information to answer the question below: The demand and supply curves are represented by the following equations: Qd=100-P and Qs=50+P . The equilibrium price is?


1
Expert's answer
2021-06-30T10:13:12-0400

"Qd = Qs"


"100-P = 50+P"


"100-50 = P + P"


"50 = 2P"


"P = \frac{50}{2}"

"P = 25"

Equilibrium price "= 25"